Which of these multiplications gives the largest answer?


(9 x 10^4) x (7 x 10^3)
(3 x 10^5) x (6 x 10^-2)
(4 x 10^4) x (3 x 10^1)
(1 x 10^5) x (4 x 10^6)
1
Expert's answer
2020-10-05T10:52:43-0400

"(1 \cdot 10^5) \cdot (4 \cdot 10^6)=4\cdot10^{11}"